

********************************************************************
UARS WEEKLY STATUS REPORT - 14 AUG 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and occulation
data during this past week. The yearly calibration using all four
deuterium lamps was completed without serious problems.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS team has continued to work towards determining a useful new operating
mode for MLS. Although there are some encouraging signs, we are not there
yet. A couple of times the scan machanism has become stuck, but we have freed
it. The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is currently off, awaiting activation on August 19, 1998.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 406 CD-ROM requests as of 14-August-1998.
There were two orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
100 days HAL (4.80) PROD (L1) (1873R-1020C-18F=835L)
61 days HAL (4.80) PROD (L3) (1873R-648C=1225L)
111 days SOL (4.62) PROD (L2-L3) (2349R-1053C=1296L)
2 days SUS (7.0) ENG (2R-2C=0L)
0 days WIN (5.10) ENG (1R-1F=1L)
The CDHF is preparing for the upgrade of the current FORTRAN compiler, version
6.3, to version 6.5. A statement of impact has been drafted and sent to the
user community this week.
The CDHF continues to identifying all patches required on the current UARS VMS
cluster operating systems for Y2K compliance. A schedule for implementing
these releases is being developed.
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5
***********
FOT
***********
REPORTING PERIOD
This report covers 01 August 1998 (Orbit 37636, GMT Day 213) through 07
August 1998 (Orbit 37740, GMT Day 219).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 213/2226 37650 Zeps Off
SUSIM 216/0105 37682 Detect. Off
MLS 217/2058 37709 Scan Test
The beta angle ranged between the angles of 62.4 degrees to 71.5 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0.0 degrees on 25 August 1998 (DOY 237).
Four (4) additional SSPP SOLSTICE Offset sequences were performed this week
on GMT days 215 through 218 at the request of SUSIM operations (and with
the agreement of SOLSTICE operations). The offsets were identical to the
regular SSPP SOLSTICE Offset sequence, which is nominally scheduled once
per week.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 31 July 1998 and 07 August 1998 are provided in
Appendix D.
The clock error ranged from 11.1 to -10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% MLS 04
% PEM 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/218 show the observatory data loss to be
20 hours, 45 minutes, 50.356 seconds (an increase of 7.872 seconds since
last report period). This is a 0.0345 percent data loss which equals a
99.9655 percent data capture for the mission.
The increase of 7.872 seconds occurred over four days:
7 0.160 seconds on Day 212 due to line errors,
7 0.160 seconds on Day 214 due to line errors,
7 7.296 seconds on Day 215 due to insufficient overlap, and
7 0.256 seconds on Day 218 due to line errors.
Some of the operational CMS workstations temporarily lost power on 07
August due to an unscheduled power outage in Building 32. This resulted in
the automatic initiation of a Rshadow disk mergeS process on the CMS. Once
the shadow merge was completed, the CMS appeared to be operating normally,
although CMS processing was extremely slow during the merge process.
Applications Processor (AP) ground system software testing has been
suspended waiting release of an operational pack containing version 10.11.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
An FOT Standard Operating Procedure is being developed outlining how to
restore files to the CMS from the incremental or full back-up tapes. The
SOP is still under review, and should be completed sometime next week.
Flight operations is investigating an alternate to sun pointing safe hold
which can provide better protection for a two battery power subsystem. The
FOT is also attempting to define a one battery scenario which would save
the spacecraft and allow a single instrument to operate or the SSPP
instruments to operate with all other instruments completely OFF.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - has been analyzed and determined
to be 2000 compatible.
User Planning System (UPS) - The UPS software will be year 2000 compatible
with Release 11 Build 3 scheduled for development in September 1998 (The
Integrator, June 1998).
Multi Satellite Operations Control Center (MSOCC) - flight operations is
Rstanding byS to provide support to this CNMOS effort. The UARS simulator
is available on request.
Command Management System (CMS) - FOT supported the CMS Year 2K test team,
as well as the CMS software maintenance team, in determining the status of
the back-up CMS system. Apparently, the login files for the various
accounts on that system are not presently set-up to support operational
type testing. The CMS system managers will be contacted as soon as
possible to get this problem rectified.
In addition, the status of the two pending CMS software releases (R7B1V1
and R7B1V2) was also investigated. It appears that both software releases
are on the back-up CMS system, but the releases must be RinstalledS before
any testing of them can be performed.
FOT met with Year 2K test team lead (Sidney Kirschner) to discuss
additional hands-on training on the CMS, as well as development of the Year
2K test plan. These meeting will probably be frequent throughout the
duration of the Year 2K testing.
Flight Dynamics Facility (FDF) - The FOT and Test Team have both requested
test planning aids from FDF, but no products have been delivered. These
products are essential to testing the CMS software.
OBC PMON FLIGHT SOFTWARE UPDATE
Flight Software - The Flight Software is complete.
Ground Support Software - Complete and delivered to the library. The
Applications Processor operational disk pack is being generated.
Verification Testing - The flight software and the test pack version of the
ground support software (Applications Processor) have been tested. Testing
of the operations pack compilation of the software will be required prior
to uplink.
IP NETWORK TESTING
Space Network (SN) - Testing of command and telemetry with the UARS
simulator completed successfully. NCC testing has been attempted but
aborted due to line configuration problems. NCC testing is on hold until
the MSOCC line configuration problem can be resolved. Performance testing
with the spacecraft will be scheduled after NCC test completion.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) - UARS emergency backup sites not scheduled at
this time. Investigating the procedure.
Ground Spaceflight Tracking and Data Network (GSTDN) - UARS emergency
backup sites not scheduled at this time. Investigating the procedure.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. Latch valve testing at Eaton is
delayed due to the unavailability of key personnel. Results are expected
in late August 1998. The Eaton Test results will be used to determine the
timing of the latch valve open times for the GSFC hardware fill test. The
command timing will be designed for execution from the UTTS by the UARS
FOT. The Propulsion Module Electronics (PME) simulation circuit design is
complete but part ordering and further work is on hold until funding issues
can be resolved.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
Question are being researched and answers provided regarding details of the
current control center software (MSOCC) and the off-line analysis software
resident on the Central Data Handling Facility (CDHF).
UARS COMMAND RE-HOST PHASE 1 DEMONSTRATION
Flight operations continues to support the update of the Lockheed Martin
SCS-21 control center system to support a phase 1 feasibility
demonstration. Requirements have been provided and questions are being
researched and answers provided for detail implementation issues. A
demonstration test script has been generated and successfully tested with
the UARS simulator.
The NCC portion of the test script was tested with the NCC 98 system on 05
August 1998. This test was complicated by line configuration problems
connecting the NCC 98 system but completed successfully. The test script
is now considered complete.
MEETINGS
The FOT attended a weekly CMS Year 2K status meeting (hosted by Carey
Myers) on 06 August 1998. Topics of discussion at the meeting were the
current status of the back-up CMS system, the status of the Year 2K testing
itself, and the timeline and deadlines related to the testing.
No UARS Battery telecon was held this week. The next telecon has been
scheduled for 25 August 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 25 August
1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 August 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37637 213/0154 98-276 OBC Clk Adj fr 20.6 to 41.2 msec/day
37650 213/2226 98-277 PEM ZEPS Power OFF
37666 215/0059 98-278 OBC Clk Adj fr 41.2 to 20.6 msec/day
37682 216/0105 98-279 SUSIM Turn off one detector
37688 216/1119 98-280 OBC Clk Adj fr 20.6 to 41.2 msec/day
37694 216/2054 98-281 MLS MLS Stop/Run
37697 217/0224 98-282 MLS MLS STOP/RUN
37709 217/2058 98-283 MLS MLS Scan Test
37739 219/2106 98-284 OBC Clk Adj fr 41.2 to 20.6 msec/day
37740 219/2337 98-285 MLS Start Nominal Scan Program
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 August 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
MLSSCANACTTMP Y-LO MLS Scan Actuator Temp 37697
MLSSCANACTTMP R-LO MLS Scan Actuator Temp 37697
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 August 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37637 213/02:01 (01 Aug 98) from 20.6 to 41.2 msec/day
37666 215/01:08 (03 Aug 98) from 41.2 to 20.6 msec/day
37688 216/11:24 (04 Aug 98) from 20.6 to 41.2 msec/day
37739 219/21:15 (07 Aug 98) from 41.2 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 07 August 1998
31 Jul 1998 - GMT Day 212 Beta = 62.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.1 0.000 0.0
2 +16.8/-22.4 +4.51/+2.92 +22.4/-16.0 26.2 1.025 19.8
3 +44.8/-11.2 +3.23/+1.97 +22.0/-15.2 26.2 1.029 19.3
07 Aug 1998 - GMT Day 219 Beta = 71.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.1 0.000 0.0
2 +11.2/ -33.6 +3.87/+1.97 +19.2/-16.0 26.7 1.034 16.8
3 +22.4/ -22.4 +2.60/+1.03 +18.4/-15.2 26.6 1.031 16.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 August 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-045 FOT See list Aug 1998 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list Jul 1998 14170 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-045 FOT See list Aug 1998 14170 Generic Late Acquisition